OpenStudy (anonymous):

Consider the function f(x)=x−15x^(1/3). a.) The domain of this function: (-inf,inf) . b.) The x-intercept(s): ? and the y-intercept(s): 0. c.) The critical number(s): ?. d.) f(x) is increasing on the interval(s) and decreasing on ? e.) f(x) has a local maximum at x =-11.18035 and a local minimum at x = 11.180344. f.) f(x) is concave up on the interval(s) and concave down on ? Can someone please help me finish this problem...

OpenStudy (cwrw238):

b) to find the x intercepts put x−15x^(1/3). = 0 and solve for x y intercepts are when x = 0 = 0 - 15*0 = 0

OpenStudy (cwrw238):

not sure what you mean by critical numbers